# DOMAIN: milDly cOnservative Model-bAsed offlINe reinforcement learning

### v2

@@ IV-B DOMAIN Optimizes a Low Bound Value
Considering the influence of sampling errors and model errors, we present the conditions for the existence of a lower bound on the value function in the DOMAIN, demonstrating that DOMAIN does not underestimate the value function for all states. Additionally, we discuss a sufficient condition under which the lower bound of DOMAIN is tighter than that of COMBO, indicating that DOMAIN exhibits weaker conservatism on offline dataset.
@@
The value function learned using Eq. (9) serves as a lower bound on the true value function, i.e. $V^{\pi}(s)\geq\widehat{V}^{\pi}(s)$, given that the following conditions are satisfied: $\int_{\mathcal{A}}\omega(s,a)\mathrm{d}a>\xi(s)\int_{\mathcal{A}}d(s,a)\mathrm{d}a$ and $\lambda\geq\delta_{l}$, where the state coefficient $\xi(s)$ and $\delta_{l}$ are defined as:
$\xi(s)=\frac{(1-f)d^{\pi_{b}}(s)\max\limits_{a}\left[\pi_{b}(a\mid s)/\pi(a\mid s)\right]+fd_{\widehat{\mathcal{M}}}^{\pi}(s)}{(1-f)d^{\pi_{b}}(s)\min\limits_{a}\left[\pi_{b}(a\mid s)/\pi(a\mid s)\right]+fd_{\widehat{\mathcal{M}}}^{\pi}(s)}\geq 1$ (15)
$\delta_{l}=\frac{(1-f)\frac{C_{r,T,\delta}R_{\max}}{(1-\gamma)\min_{s,a}\{\sqrt{|\mathcal{D}|}\}}}{\min_{s}\left\{\frac{(\xi(s)-1){\color[rgb]{0,0,0}\int_{\mathcal{A}}d(s,a)\mathrm{d}a}}{\left[(1-f)d^{\pi_{b}}(s)\max_{a}\left\{\frac{\pi_{b}(a\mid s)}{\pi(a\mid s)}\right\}+fd_{\widehat{\mathcal{M}}}^{\pi}(s)\right]}\right\}}$
$+\frac{f[\max\limits_{s,a}\left\{\left|r_{\mathcal{M}}-r_{\mathcal{M}}\right|\right\}+\frac{2\gamma R_{\max}}{1-\gamma}\max\limits_{s,a}\left\{D_{TV}\left(T_{\mathcal{M}},T_{\widehat{\mathcal{M}}}\right)\right\}]}{\min_{s}\left\{\frac{(\xi(s)-1){\color[rgb]{0,0,0}\int_{\mathcal{A}}d(s,a)\mathrm{d}a}}{\left[(1-f)d^{\pi_{b}}(s)\max_{a}\left\{\frac{\pi_{b}(a\mid s)}{\pi(a\mid s)}\right\}+fd_{\widehat{\mathcal{M}}}^{\pi}(s)\right]}\right\}}$
@@
Furthermore, DOMAIN does not underestimate the value function for all states $s$. It can be inferred that the value function is underestimated only when the condition in Eq. (15) is satisfied. However, there is a possibility of overestimating the data with high probabilities in the distribution $d(s,a)$ of the offline data set. Since the data points $(s,a)$ with a higher frequency of occurrence in region $\mathcal{D}_{\text{offline }}$ are considered as in-distribution data, the learned environment model $\widehat{T}_{\phi}$ exhibits smaller prediction errors for such data, resulting in smaller values of $\omega(s,a)$. As a result, it becomes challenging to satisfy the condition $\int_{\mathcal{A}}\omega(s,a)\mathrm{d}a>\xi(s)\int_{\mathcal{A}}d(s,a)\mathrm{d}a$, leading to overestimation. This aligns with the adaptive adjustment of penalties for reward.
@@ IV-C DOMAIN Guarantees the Safety of Policy Improvement
Let $\pi^{*}(a\mid s)$ be the policy optimized by DOMAIN. Then, $\pi^{*}(a\mid s)$ represents a $\zeta$safe policy improvement over $\pi_{b}(a\mid s)$ in the actual MDP $\mathcal{M}$, satisfying $J\left(\mathcal{M},\pi^{*}\right)-J\left(\mathcal{M},\pi_{b}\right)\geq\zeta$ with high probability $1-\delta$. The value of $\zeta$ is given by:
@@ V Experiments
This section aims to address the following questions: 1) How does the DOMAIN method compare to state-of-the-art RL algorithms in standard offline RL benchmarks? 2) Can the DOMAIN method handle tasks that require generalization to OOD behaviors?
Figure 2: Description of Gym-Mujoco and Maze2D Tasks. The Gym-Mujoco benchmark includes Halfcheetah, Hopper and Walker2d tasks. The Maze2D benchmark includes three maze layout tasks.
@@ V-A The Implementation Details of Experiments
Experimental Datasets. We employ Gym-Mujoco and Maze2D benchmarks to test the performance of DOMAIN. The Gym-Mujoco tasks, including Halfcheetah-v2, Hopper-v2, and Walker2d-v2, are commonly studied in OpenAI Gym [35], simulated using the MuJoCo physics engine. The Maze2D domain, including Umaze-v1, Medium-v1 and Large-v1 three maze layouts, is a navigation task requiring the agent to reach a fixed goal location.
Fig. 2 depicts the Halfcheetah, Hopper, Walker2d and Maze2D four tasks. For three Gym-Mujoco tasks, the goal is to control the robot’s various joints to achieve faster and more stable locomotion. By applying RL methods to robot control, the agent observes the state of the MuJoCo environment and controls the robot’s joints based on its learned policy. For Maze2D tasks, the goal is to find the shortest path to reach the fixed goal location. The offline datasets of Gym-Mujoco and Maze2D benchmarks are based on the D4RL dataset [36].
For three Gym-Mujoco tasks, three types of recorded datasets are considered: Medium, Medium-replay, and Medium-expert. The “Medium” dataset is generated by collecting 1 million samples from a policy that undergoes early stopping during online training using SAC [27]. The “Medium-replay” dataset records all samples during the training process, reaching a medium level. The “Medium-expert” dataset is a combination of expert and sub-optimal data [36].
@@
1) Model Training: The environment model is approximated using a probabilistic network that outputs a Gaussian distribution to obtain the next state and reward. Similar to the approach in [19], we set $N$ as 7 and select the best 5 models. Each model consists of a 4-layer feedforward neural network with 200 hidden units. The model training employs maximum likelihood estimation with a learning rate of 1e-4 and Adam optimizer. The 256 data samples are randomly drawn from the offline dataset $\mathcal{D}_{\text{offline }}$ to calculate the maximum likelihood term.
@@
3) Policy Optimization: The training step is set as $1$ million and half a million for Gym-Mujoco tasks and Mazed2D tasks respectively. Policy optimization is based on the SAC framework. The critic network $Q_{\omega}$ and the policy network $\pi_{\theta}$ adopt a 2-layer feedforward neural network with $256$ hidden units. The learning rates for them are set to $43e-$ 4 and $41e-$ 4 respectively. The ratio of model data in the dataset is $f=0.5$. The entropy regularization coefficient $\alpha$ in Eq. (5) is automatically adjusted, with the entropy target set to -dim(A), and the learning rate for the self-tuning network is set to $41e-$ 4.
@@
In practical policy optimization, the obtained adaptive sampling distribution is utilized to select data with high uncertainty and penalize their $Q$-values. Following [13, 21], we sample 10 actions in every high uncertainty state based on the uniform policy Unif(a) and the learned policy $\pi(a\mid s)$, and then estimate the log-sum-exp term in (9) using importance sampling.
Parameter Tuning. The quality of model data is influenced by the environment model $\widehat{T}_{\phi}$, the rollout policy $\mu(\cdot\mid s)$, and the rollout horizon $H$. The weight coefficient $\lambda$ of the regularization term significantly affects the stability of agent training. The rollout policy for Gym-Mujoco and Mazed2D benchmarks are the current learned policy and the uniform policy, respectively. Table III and Table VI give the results of Gym-Mujoco tasks and Maze2D tasks for different parameters, respectively. The detailed parameter tuning results in different tasks are as follows.
1) Rollout horizon $H$ : For Gym-Mujoco tasks, a short-horizon model rollout is performed [19, 24]. The rollout horizon $H$ is chosen from the set $\{1,5\}$. We ues $H=1$ for Walker2d-v2 tasks, and $H=5$ for Hopper-v2 and Halfcheetah-v2. For Maze2D tasks, the rollout horizon $H$ is chosen from the set $\{5,10,100\}$. We ues $H=5$ for Maze2D-umaze task, $H=10$ for Maze2D-medium task, and $H=100$ for Maze2D-large task.
2) Weight coefficient $\lambda$ : The parameter $\lambda$ reflects different strengths of regularization. For all tasks, the parameter $\lambda$ is chosen from the set $\{0.5,5\}$. A larger $\lambda$ is needed in narrower data distribution. A smaller is needed in diverse data distribution. For three Mazed2D tasks, Gym-Mujoco medium-replay datasets and halfcheetah-medium datasets, $\lambda$ is chosen as 0.5, while for the remaining datasets, $\lambda$ is chosen as 5.
TABLE II: Results for the Gym-Mujoco tasks during the last 5 iterations of training averaged over 5 seeds. $\pm$ captures the standard deviation over seeds. Bold indicates the highest score. Highlighted numbers indicate results within 2% of the most performance algorithm.
@@ V-B Experimental Results
Results on the Gym-Mujoco tasks. To answer question 1, this section compares recent model-free algorithms, such as MOAC [40], IQL [37], and GORL [41], as well as model-based algorithms, including ARMOR [25], MAPLE [39], MOReL [20], COMBO [21], and RAMBO [24]. The performance of these algorithms is evaluated based on the cumulative return in the robot motion tasks within the MuJoCo environment. A higher score indicates better performance in terms of more stable and faster locomotion control. For comparison, the scores are normalized between 0 (random policy score) and 100 (expert policy score) [36]. The normalized score $\tilde{S}$ is computed by:
@@
Table II presents the normalized scores of different methods. The reported scores are the normalized scores of the learned policies during the last 5 iterations, averaged over 5 random seeds. Note that the average of the last 1 [21, 25], 5 and 10 [24] iterations are usually selected as final results. The number of the last iteration of different algorithms should be the same for fair comparison. For our algorithm, the results are basically consistent in these three cases. Here we choose the average of the last 5 iterations as the final result. The results for all methods are obtained from the original papers.
The table shows that the DOMAIN method achieves the best performance in three out of the nine datasets, comparable results in three out of the remaining six settings, and slightly poor performance in the three medium settings. It is difficult for model-based offline RL algorithms to perform well on all datasets due to the differences between various datasets. Therefore, most offline RL algorithms show the advantages of the algorithm by comparing the average values under different tasks. Overall, compared to other RL algorithms, the DOMAIN algorithm achieves the best average performance across multiple datasets. Corresponding learning curves are shown in Fig. 5.
Moreover, Table II shows that the performance of the “Medium-expert” dataset in three tasks is almost always superior to that of “Medium” and “Medium-replay” datasets for any algorithms. The difference between “Medium” and “Medium-replay” is that “Medium-replay” contains more data at different levels, which can facilitate agent exploration. However, the algorithm may not necessarily perform better than “Medium” under “Medium-replay” for different tasks. This mainly depends on the adaptability of the algorithm itself to diverse data and the coverage of the data.
TABLE IV: Results for tasks that require generalization during the last 5 iterations of training.
Environment DOMAIN COMBO MoReL MOPO MBPO CQL SAC BRAC-p BEAR
Halfcheetah-jump $\mathbf{93.65\pm 3.02}$ $45.02$ $28.26$ $34.61$ $26.19$ $8.23$ $-26.64$ $10.87$ $2.39$
TABLE V: Results for the Maze2D tasks during the last 5 iterations of training averaged over 5 seeds. $\pm$ captures the standard deviation over seeds. Bold indicates the highest score.
Type Environment Ours Model-based baseline Model-free baseline
DOMAIN ROMI-BCQ COMBO MoReL MOPO BEAR CQL BRAC BAIL
Sparse Maze2D-umaze $\mathbf{142.6\pm 7.7}$ $139.5$ $76.4$ $-14.4$ $-14.8$ $65.7$ $18.9$ $-9.2$ $44.7$
Maze2D-medium $\mathbf{154.4\pm 12.0}$ $82.4$ $68.5$ $-4.8$ $37.4$ $25.0$ $14.6$ $70.0$ $19.7$
Maze2D-large $\mathbf{152.1\pm 42.8}$ $83.1$ $14.1$ $-2.3$ $-0.8$ $81.0$ $16.0$ $0.2$ $4.4$
Average Score $\mathbf{149.7\pm 20.8}$ $101.7$ $53.0$ $-7.2$ $7.3$ $57.2$ $16.5$ $20.3$ $22.9$
TABLE VI: The results of Maze2D tasks for different parameters. Bolded and highlighted numbers indicate the best performance
Type Environment $H=5$ $H=10$ $H=50$ $H=100$
Sparse Maze2D-umaze $\mathbf{142.63}$ $66.94$ $74.16$ $-15.82$
Maze2D-medium $107.03$ $\mathbf{154.39}$ $126.33$ $114.15$
Maze2D-large $52.48$ $112.08$ $107.57$ $\mathbf{152.14}$
[…]
Figure 6: Corresponding learning curves for Table V. The shaded area represents the standard variance of the return. The shaded area represents the standard variance of the return.
The datasets used in both cases are collected by training the SAC algorithm for 1 million steps and are similar to the “Medium-expert” dataset in D4RL. Therefore, this dataset serves as the foundation dataset, and the rewards are appropriately modified to obtain the final dataset for training. Table IV presents the normalized scores of different algorithms in the halfcheetah-jump environment. The results for model-based algorithms are sourced from the work of [21], while the results for model-free algorithms (BEAR [9], BRAC [11], SAC [27], and CQL [13]) are obtained from [19]. It can be observed that the DOMAIN algorithm performs the best, demonstrating stronger generalization capabilities and robustness.
Results on the Maze2D tasks. The Maze2D tasks include three different maze layouts. The names of three tasks in the D4RL dataset are “Maze2d-umaze-sparse”, “Maze2d-medium-sparse” and “Maze2d-large-sparse”, respectively. These tasks are challenging for RL methods since the reward is sparse. The score calculation of Maze2D is the same as the Gym-Mujoco tasks. The random policy scores under Umaze, Medium, and Large mazed layouts are $23.85$, $13.13$, and $6.7$ respectively. The expert policy scores under Umaze, Medium, and Large mazed layouts are $161.86$, $277.39$, and $273.99$ respectively.
Table V compares the normalized scores of DOMAIN with that of other current offline RL algorithms on Maze2D tasks, where model-based algorithms include ROMI [42], MOReL [20], COMBO [21] and MOPO [19], and model-free algorithms include BEAR [9], BRAC [11], BAIL [43] and CQL [13]. Apart from ROMI, the original papers of the above methods don’t give the results on the Maze2D benchmark. Therefore, the results of the above method are sourced from the work of [42]. Table V shows that the performance of DOMAIN is much better than other methods, which also shows DOMAIN has advantages in sparse reward tasks. Corresponding learning curves are shown in Fig. 6.
Moreover, the methods DOMAIN and COMBO, which regularize the model data to incorporate conservatism, perform better than the methods MOPO and MOReL that use model uncertainty as a penalty for rewards on the Maze2D tasks. This shows the unreliability of estimating model uncertainty. DOMAIN performs much better than COMBO, which further verifies the advantage of the adaptive sampling distribution. The maximum rollout horizons of Umaze, Medium and Large are $300$, $600$ and $800$ respectively. The tasks are becoming more and more complex. The parameter tuning of rollout length in Table VI indicates that the large rollout length is beneficial for the performance of long-horizon tasks.
TABLE VII: The parameter configuration of our designed simple MDP.
@@ V-C Visualisation of DOMAIN Performance
Due to the high dimension of states and actions in the D4RL tasks, visualizing the performance is greatly challenging. Therefore, a simple example MDP is defined for illustrative purposes following [24]. For this MDP, both the state space $\mathcal{S}$ and action space $\mathcal{A}\in[-1,1]$ are one-dimensional. The reward is defined as $R(s,a)=s$, where a larger value of state $s$ leads to a higher reward. The dynamic transition model is given by $T(s^{\prime}\mid s,a)=\mathcal{N}(\mu,\sigma)$ with the parameters shown in Table VII, indicating that the next state depends on the current action, independent of the initial state. To visualize the DOMAIN exploration in OOD region, we sample actions from $\mathcal{N}(\mu=0,\sigma=0.35)$, which means that the behavioral policy for the offline dataset is random policy.
@@
TABLE VIII: The analysis results for parameter $H$ and $\lambda$. Bold indicates the highest score in ablation experiments. Highlighted numbers indicate the optimal results in Table III.

### v3

@@ IV-B DOMAIN Optimizes a Low Bound Value
Considering the influence of sampling errors and model errors, we present the conditions for the existence of a lower bound on the value function in the DOMAIN, and demonstrate that DOMAIN does not underestimate the value function for all states. Additionally, we discuss a sufficient condition under which the lower bound of DOMAIN is tighter than that of COMBO, indicating that DOMAIN exhibits weaker conservatism on offline dataset.
@@
The value function learned using Eq. (9) serves as a lower bound on the true value function, i.e. $V^{\pi}(s)\geq\widehat{V}^{\pi}(s)$, given that the following conditions are satisfied: $\sum_{a}\omega(s,a)>\xi(s)\sum_{a}d(s,a)$ and $\lambda\geq\delta_{l}$, where the state coefficient $\xi(s)$ and $\delta_{l}$ are defined as:
$\xi(s)=\frac{(1-f)d^{\pi_{b}}(s)\max\limits_{a}\left[\pi_{b}(a\mid s)/\pi(a\mid s)\right]+fd_{\widehat{\mathcal{M}}}^{\pi}(s)}{(1-f)d^{\pi_{b}}(s)\min\limits_{a}\left[\pi_{b}(a\mid s)/\pi(a\mid s)\right]+fd_{\widehat{\mathcal{M}}}^{\pi}(s)}\geq 1$
$\delta_{l}=\frac{(1-f)\frac{C_{r,T,\delta}R_{\max}}{(1-\gamma)\min_{s,a}\{\sqrt{|\mathcal{D}|}\}}}{\min_{s}\left\{\frac{(\xi(s)-1)\sum_{a}d(s,a)}{\left[(1-f)d^{\pi_{b}}(s)\max_{a}\left\{\frac{\pi_{b}(a\mid s)}{\pi(a\mid s)}\right\}+fd_{\widehat{\mathcal{M}}}^{\pi}(s)\right]}\right\}}$ (15)
$+\frac{f[\max\limits_{s,a}\left\{\left|r_{\mathcal{M}}-r_{\mathcal{M}}\right|\right\}+\frac{2\gamma R_{\max}}{1-\gamma}\max\limits_{s,a}\left\{D_{TV}\left(T_{\mathcal{M}},T_{\widehat{\mathcal{M}}}\right)\right\}]}{\min_{s}\left\{\frac{(\xi(s)-1)\sum_{a}d(s,a)}{\left[(1-f)d^{\pi_{b}}(s)\max_{a}\left\{\frac{\pi_{b}(a\mid s)}{\pi(a\mid s)}\right\}+fd_{\widehat{\mathcal{M}}}^{\pi}(s)\right]}\right\}}$
@@
Furthermore, DOMAIN does not underestimate the value function for all states $s$. It can be inferred that the value function is underestimated only when the condition in Eq. (15) is satisfied. However, there is a possibility of overestimating the data with high probabilities in the distribution $d(s,a)$ of the offline data set. Since the data points $(s,a)$ with a higher frequency of occurrence in region $\mathcal{D}_{\text{offline }}$ are considered as in-distribution data, the learned environment model $\widehat{T}_{\phi}$ exhibits smaller prediction errors for such data, resulting in smaller values of $\omega(s,a)$. As a result, it becomes challenging to satisfy the condition $\sum_{a}\omega(s,a)>\xi(s)\sum_{a}d(s,a)$, leading to overestimation. This aligns with the adaptive adjustment of penalties for reward.
@@ IV-C DOMAIN Guarantees the Safety of Policy Improvement
Let $\pi^{*}(a\mid s)$ be the policy optimized by DOMAIN. Then, $\pi^{*}(a\mid s)$ represents a $\zeta$-safe policy improvement over $\pi_{b}(a\mid s)$ in the actual MDP $\mathcal{M}$, satisfying $J\left(\mathcal{M},\pi^{*}\right)-J\left(\mathcal{M},\pi_{b}\right)\geq\zeta$ with high probability $1-\delta$. The value of $\zeta$ is given by:
@@
Figure 2: The description of Mujoco tasks, including Halfcheetah, Hopper and Walker2d tasks.
@@ V Experiments
This section aims to evaluate the performance of DOMAIN on standard benchmark, and gives the details of experimental setup, results and parameter analysis.
@@ V-A The Implementation Details of Experiments
Experimental Datasets. We employ Gym-Mujoco benchmark to test the performance of DOMAIN. The Gym-Mujoco tasks, including Halfcheetah-v2, Hopper-v2, and Walker2d-v2, are commonly studied in OpenAI Gym [35], simulated using the MuJoCo physics engine. Fig. 2 depicts the above tasks. The goal is to control the robot’s various joints to achieve faster and more stable locomotion. [36].
For three Mujoco tasks, three types of recorded datasets are considered: Medium, Medium-replay, and Medium-expert. The “Medium” dataset is generated by collecting 1 million samples from a policy that undergoes early stopping during online training using SAC [27]. The “Medium-replay” dataset records all samples during the training process, reaching a medium level. The “Medium-expert” dataset is a combination of expert and sub-optimal data [36].
@@
1) Model Training: The environment model is approximated by a probabilistic network that outputs a Gaussian distribution to obtain the next state and reward. We set $N$ as 7 and select the best 5 models. Each model consists of a 4-layer feedforward network with $200$ hidden units. The model employs maximum likelihood estimation with a learning rate of $41e-$ 4 and Adam optimizer. The $256$ data samples are randomly drawn from $\mathcal{D}_{\text{offline }}$ to train model.
@@
3) Policy Optimization: Policy optimization is based on the SAC framework. The critic network $Q_{\omega}$ and the policy network $\pi_{\theta}$ adopt a 2-layer feedforward network with $256$ hidden units. The learning rates for them are set to $43e-$ 4 and $41e-$ 4 respectively. The ratio of model data is $f=0.5$. The entropy regularization coefficient $\alpha$ in Eq. (5) is automatically adjusted, with the entropy target set to -dim(A), and the learning rate for the self-tuning network is set to $41e-$ 4.
@@
In practical policy optimization, the obtained adaptive sampling distribution is utilized to select model data with high uncertainty and penalize their $Q$-values. Following [13, 21], we sample 10 actions in every high uncertainty state based on the uniform policy $Unif(a)$ and the learned policy $\pi(a\mid s)$, and then estimate the log-sum-exp term in (9) using importance sampling.
Parameter Tuning. The quality of model data is influenced by the environment model $\widehat{T}_{\phi}$, the rollout policy $\mu(\cdot\mid s)$, and the rollout horizon $H$. The weight coefficient $\lambda$ of the regularization term significantly affects the stability of agent training. The rollout policy is the current learned policy. Table III gives the results of Gym-Mujoco tasks for different parameters. The detailed parameter tuning results in different tasks are as follows.
1) Rollout horizon $H$ : The rollout horizon $H$ is chosen from the set $\{1,5\}$. We ues $H=1$ for Walker2d-v2 tasks, and $H=5$ for Hopper-v2 and Halfcheetah-v2.
2) Weight coefficient $\lambda$ : The parameter $\lambda$ reflects different strengths of regularization. The parameter $\lambda$ is chosen from the set $\{0.5,5\}$. For three medium-replay datasets and halfcheetah-medium datasets, $\lambda$ is chosen as 0.5, while for the remaining datasets, $\lambda$ is chosen as 5.
TABLE II: Results for D4RL tasks during the last 5 iterations of training averaged over 5 seeds. $\pm$ captures the standard deviation. Bold indicates the highest score. Highlighted numbers indicate results within 2% of the most performance algorithm.
@@ V-B Experimental Results
Results on the D4RL tasks. This section compares recent model-free algorithms, such as MOAC [40], IQL [37], and GORL [41], as well as model-based algorithms, including ARMOR [25], MAPLE [39], MOReL [20], COMBO [21], and RAMBO [24]. The performance of these algorithms is evaluated based on the cumulative return in the robot motion tasks within the MuJoCo environment. A higher score indicates better performance in terms of more stable and faster locomotion control. For comparison, the scores are normalized between 0 (random policy score) and 100 (expert policy score) [36]. The normalized score $\tilde{S}$ is computed by:
@@
Table II presents the normalized scores of different methods. The reported scores are the normalized scores of the learned policies during the last 5 iterations, averaged over 5 random seeds. The results for all methods are obtained from the original papers. The table shows that the DOMAIN method achieves the best performance in three out of the nine datasets, comparable results in three out of the remaining six settings, and slightly poor performance in the three medium settings. Overall, DOMAIN can achieve better average performance across multiple datasets. Corresponding learning curves are shown in Fig. 5.
TABLE IV: The parameter configuration of our designed simple MDP.
@@ V-C Visualisation of DOMAIN Performance
Due to the high dimension of states and actions in the D4RL tasks, visualizing the performance is greatly challenging. Therefore, a simple example MDP is defined for illustrative purposes following [24]. For this MDP, both the state space $\mathcal{S}$ and action space $\mathcal{A}\in[-1,1]$ are one-dimensional. The reward is defined as $R(s,a)=s$, where a larger value of state $s$ leads to a higher reward. The dynamic transition model is given by $T(s^{\prime}\mid s,a)=\mathcal{N}(\mu,\sigma)$ with the parameters shown in Table IV, indicating that the next state depends on the current action, independent of the initial state. To visualize the DOMAIN exploration in OOD region, we sample actions from $\mathcal{N}(\mu=0,\sigma=0.35)$, which means that the behavioral policy for the offline dataset is random policy.
@@
TABLE V: The analysis results for parameter $H$ and $\lambda$. Bold indicates the highest score in ablation experiments. Highlighted numbers indicate the optimal results in Table III.